OpenStudy (hari5719):

calculate the no of moles in (a) 24 g of neon gas

OpenStudy (hari5719):

@nincompoop

OpenStudy (nincompoop):

\[24g_{Ne} \times \frac{ 1mole_{Ne} }{ 20.1797g_{Ne} } =\]

OpenStudy (nincompoop):

turning a rabbit to frog \[Rabbit \times \frac{ Frog }{ Rabbit }= Frog \] turning a frog to rabbit \[Frog \times \frac{ Rabbit }{ Frog }=Rabbit\] if you haven't taken first-level algebra, then you are not equipped to take this class.